# Removing neural signal artifacts with autoencoder-targeted adversarial transformers (AT-AT)

### v1

@@ 1 Introduction
Advanced brain-computer interfaces (BCIs) promise to unlock exciting possibilities in both rehabilitation (Huang et al., 2022) and augmentation (Cinel et al., 2019). Specific potential applications include neuroprosthetic devices (Choi and Liu, 2025), nonverbal communication systems (Kashihara, 2014), paralysis aids (Chaudhary et al., 2021), and beyond. However, poor signal quality—particularly for noninvasive brain-computer interfaces that attempt to avoid intracranial implantation processes (Sun et al., 2018)—presents a major barrier to deployment. In this paper, we seek to improve the signal quality of non-invasive techniques like electroencephalography (EEG) by eliminating the primary threat of electromyographic (EMG) interference (Liang et al., 2021). A more comprehensive examination of this issue, alongside end-use applications and geometric machine learning methodologies, is provided in Choi (2025).
Conventional algorithmic solutions for filtering EMG interference in raw EEG largely center upon linear signal processing methods like canonical correlation analysis (CCA) and independent component analysis (ICA) (Krim and Viberg, 1996; Iriarte et al., 2003; De Clercq et al., 2006), which can be inadequate in elaborate single-channel source separation scenarios (Zhang et al., 2021a). The underlying hypothesis guiding much of the work presented herein—in the lineage of prior work (Choi et al., 2025)—is that ML-based signal processing algorithms possess the requisite expressivity to potentially handle complex single-channel interference cases. Yet while recent progress in the field has demonstrated promise in terms of ML-based signal processing (Zhang et al., 2021b; Yin et al., 2023; Cui et al., 2024), issues remain in terms of expensive model footprints, latency issues, and undesirable levels of signal degradation. As emphasized in prior art (Wang et al., 2020; Pacini et al., 2024), large model footprints can be prohibitive for many real-world, resource-constrained BCI applications.
While CNN-based EEG-EMG signal filtration algorithms take advantage of the spatial locality inherent in EEG processing (Zhang et al., 2021b), transformers can harness both the spatial locality and non-locality relevant to neural signal reconstruction (Vaswani et al., 2017; Cho et al., 2023). Prior literature exists on the application of transformers to EEG data (Pu et al., 2022; Wang et al., 2024); on the whole, however, transformer denoising results remain suboptimal compared to the state-of-the-art (SOTA) CNN-based approach (Zhang et al., 2021b). We anticipate that the development of more effective transformer application site determination methods—as well as non brute-force approaches to tokenization and target site optimization—may enable superior downstream performance; these ideas have yet to be rigorously considered in EEG transformer literature.
@@
    To address point 1, we present the marriage of a signal-processing transformer with the principles of a generative adversarial network (Creswell et al., 2018). By training the generative transformer in an adversarial environment, the network output can be more carefully regulated—and thus made to precisely mimic that of pristine neural signal (leading to superior reconstructive accuracy). While adversarial transformers have been demonstrated in disjoint generative applications (Hudson and Zitnick, 2021; Kasem et al., 2019), they have yet to be applied to neural blind source separation.
@@
Figure 1: The autoencoder-targeted adversarial transformer (AT-AT) system architecture. The initial signal is passed to the LSTM-CNN (LC) targeting model and denoising autoencoder (upper panels); the transformer is subsequently invoked on select target sites (lower diagram). Design incorporates elements featured in Choi (2025); Choi et al. (2025).
The subsequent autoencoder-targeted adversarial transformer system (AT-AT) will be employed—as a lightweight, autoencoder-connected solution—to remediate the primary noise barrier inhibiting effective neural interfaces. In this study, we build an AT-AT model and validate AT-AT’s denoising performance in a standardized semi-synthetic setting on a major benchmark blind source separation (BSS) dataset. Though outside the primary scope of this study, the Appendix also provides a preliminary validation of AT-AT performance in a real-world setting (with no “ground truth" signal present); this real-world application (in conjunction with a manifold learning pipeline) is formally covered in depth in Choi (2025).
@@ 2 Methods
In order to validate the proposed AT-AT denoising system, we employed the benchmark EEGdenoiseNet dataset (Zhang et al., 2021a). While other available datasets were surveyed (Koelstra et al., 2011; Van Veen et al., 2019; Kaya et al., 2018), EEGdenoiseNet has emerged in the BCI field as a major benchmark for training and validating published ML denoising algorithms (Zhang et al., 2021a, b; Yin et al., 2023; Cui et al., 2024), thereby enabling objective comparison of AT-AT against current gold standards. Moreover, EEGdenoiseNet surpasses alternatives in terms of both overall scope and the ability for ground truth verification—crucial for benchmark testing. We conduct a preliminary validation of AT-AT in line with prior autoencoder-based work (Choi et al., 2025), using delineated methods from Zhang et al. (2021a) to construct a semi-synthetic contamination problem. We built and trained the AT-AT system (Figure 1) on two train mixtures and two test mixtures at -7 dB and 2 dB with an EEG-EMG BSS model objective. As alluded to in the Introduction, a major engineering goal of the AT-AT system was to demonstrate a targeted application of attention networks for filtration with a constrained model footprint and training time. However, to appropriately assess model training time in a live BSS context (as previously noted in Choi et al. (2025)), one must consider not only model training time but also the total combined duration of EEG and EMG segments used in a single training run. In other words, the model training time is not merely the training length itself, but rather the maximum of the training length and the duration of neural data needed for a complete training run. In light of this, for the AT-AT system, we restricted training data to 120 EEG segments contaminated with high-variance EMG artifacts totaling 240s in combined duration. Thus, four minutes’ worth of data are required to support the AT-AT training runs described in this study.
@@
In accordance with methods for inferring signal SNR before running a subsequent tailored denoising model, we first run an initial upstream model based on a hybrid LSTM-CNN (LC) architecture to determine the appropriate SNR target level (Choi et al., 2025); details of this upstream model are included in the Appendix. Subsequently, as a first denoising pass, the AT-AT architecture employs a convolutional autoencoder with batch normalization; this architecture showed promise for EEG-EMG denoising in prior work (Choi et al., 2025). This autoencoder structure performs reasonably well in lower-noise settings but still demonstrates suboptimal performance on especially noisy reconstruction sites. Thus, to rectify the shortcomings of this convolutional autoencoder, we selectively invoke a time-series transformer model to handle more intensive instances of reconstruction. Introducing a second model architecture comes at the cost of a more complex parameter space search and longer training run, so the time-series transformer is only used when necessary in order to accommodate a more intensive retraining process. This novel hybrid use of autoencoders and transformers in conjunction is a key component of the AT-AT system; an evaluation of a sans-transformer method is outside of the scope of this study, but we refer readers to Choi et al. (2025).
@@
One key challenge stemming from this combination of autoencoder and transformer architectures was that of spectral preservation. The selective insertion of time-series transformer-generated signals can introduce irregularities into the model output and prevent the model output characteristics from appropriately resembling the spectral characteristics of authentic EEG data. To achieve a well-regulated output compliant with the physical properties of EEG, we implemented an adversarial training component to our autoencoder-transformer hybrid architecture. Specifically, as mentioned in the Introduction, we treated the transformer-augmented autoencoder output as the “generator" and implemented a convolutional neural network-based (1D-CNN-based) discriminator model to distinguish between authentic versus model-reconstructed EEG. This generator-discriminator pair was folded into a five-cycle GAN training loop; ablation-wise, in line with prior work (Choi et al., 2025), we found that the addition of adversarial training led to $>$10% gains in spectral filtration accuracy.
After building and validating the AT-AT system across a partitioned 100-sample test set on the benchmark semi-synthetic contamination problem, we also implemented additional end-use case validation. In the EEGdenoiseNet context, performance can be assessed via direct comparison with the synthetically contaminated ground truth samples. But in real-world environments, the only way to assess the performance of our signal processing pipeline is by measuring improvement on downstream classification tasks. Thus, in order to demonstrate the potential of an AT-AT model at unlocking better BCI outcomes, we deployed AT-AT onto a notable BCI use case: digit- versus non-digit classification on the MindBigData EEG dataset (Vivancos and Cuesta, 2022) (see Appendix for further details). This downstream application ultimately incorporated an extensive geometric machine learning pipeline which is discussed in Choi (2025) (but falls outside of the scope of this paper).
@@ 3 Results & Discussion
Figure 3: AT-AT performance relative to major deep learning benchmarks. [A], [B], and [C] refer to Zhang et al. (2021b); Yin et al. (2023) and Cui et al. (2024), respectively. All values are best extrapolated from reported documentation (best-in-class values in bold).
@@
Examining our system on real-world use cases is an additional key component of AT-AT validation. As discussed in the Appendix and formally treated in Choi (2025), we implemented AT-AT in conjunction with an extensive downstream manifold learning pipeline on the real-world imagined digit- versus non-digit classification problem. The results of this use case deployment (which are informally included in the Appendix) offer early evidence that AT-AT may confer utility in real-world BCI settings.

### v2

@@ 1 Introduction
Advanced brain-computer interfaces (BCIs) promise to unlock exciting possibilities in both rehabilitation [10] and augmentation [6]. Specific potential applications include neuroprosthetic devices [4], nonverbal communication systems [14], paralysis aids [1], and beyond. However, poor signal quality—particularly for noninvasive brain-computer interfaces that attempt to avoid intracranial implantation processes [22]—presents a major barrier to deployment. In this paper, we seek to improve the signal quality of non-invasive techniques like electroencephalography (EEG) by eliminating the primary threat of electromyographic (EMG) interference [18]. A more comprehensive examination of this issue, alongside end-use applications and geometric machine learning methodologies, is provided in [3].
Conventional algorithmic solutions for filtering EMG interference in raw EEG largely center upon linear signal processing methods like canonical correlation analysis (CCA) and independent component analysis (ICA) [17, 12, 9], which can be inadequate in elaborate single-channel source separation scenarios [31]. The underlying hypothesis guiding much of the work presented herein—in the lineage of prior work [5]—is that ML-based signal processing algorithms possess the requisite expressivity to potentially handle complex single-channel interference cases. Yet while recent progress in the field has demonstrated promise in terms of ML-based signal processing [32, 30, 8], issues remain in terms of expensive model footprints, latency issues, and undesirable levels of signal degradation. As emphasized in prior art [28, 19], large model footprints can be prohibitive for many real-world, resource-constrained BCI applications.
While CNN-based EEG-EMG signal filtration algorithms take advantage of the spatial locality inherent in EEG processing [32], transformers can harness both the spatial locality and non-locality relevant to neural signal reconstruction [25, 2]. Prior literature exists on the application of transformers to EEG data [20, 27]; on the whole, however, transformer denoising results remain suboptimal compared to the state-of-the-art (SOTA) CNN-based approach [32]. We anticipate that the development of more effective transformer application site determination methods—as well as non brute-force approaches to tokenization and target site optimization—may enable superior downstream performance; these ideas have yet to be rigorously considered in EEG transformer literature.
@@
    To address point 1, we present the marriage of a signal-processing transformer with the principles of a generative adversarial network [7]. By training the generative transformer in an adversarial environment, the network output can be more carefully regulated—and thus made to precisely mimic that of pristine neural signal (leading to superior reconstructive accuracy). While adversarial transformers have been demonstrated in disjoint generative applications [11, 13], they have yet to be applied to neural blind source separation.
@@
Figure 1: The autoencoder-targeted adversarial transformer (AT-AT) system architecture. The initial signal is passed to the LSTM-CNN (LC) targeting model and denoising autoencoder (upper panels); the transformer is subsequently invoked on select target sites (lower diagram). Design incorporates elements featured in [3, 5].
The subsequent autoencoder-targeted adversarial transformer system (AT-AT) will be employed—as a lightweight, autoencoder-connected solution—to remediate the primary noise barrier inhibiting effective neural interfaces. In this study, we build an AT-AT model and validate AT-AT’s denoising performance in a standardized semi-synthetic setting on a major benchmark blind source separation (BSS) dataset. Though outside the primary scope of this study, the Appendix also provides a preliminary validation of AT-AT performance in a real-world setting (with no “ground truth" signal present); this real-world application (in conjunction with a manifold learning pipeline) is formally covered in depth in [3].
@@ 2 Methods
In order to validate the proposed AT-AT denoising system, we employed the benchmark EEGdenoiseNet dataset [31]. While other available datasets were surveyed [16, 24, 15], EEGdenoiseNet has emerged in the BCI field as a major benchmark for training and validating published ML denoising algorithms [31, 32, 30, 8], thereby enabling objective comparison of AT-AT against current gold standards. Moreover, EEGdenoiseNet surpasses alternatives in terms of both overall scope and the ability for ground truth verification—crucial for benchmark testing. We conduct a preliminary validation of AT-AT in line with prior autoencoder-based work [5], using delineated methods from [31] to construct a semi-synthetic contamination problem. We built and trained the AT-AT system (Figure 1) on two train mixtures and two test mixtures at -7 dB and 2 dB with an EEG-EMG BSS model objective. As alluded to in the Introduction, a major engineering goal of the AT-AT system was to demonstrate a targeted application of attention networks for filtration with a constrained model footprint and training time. However, to appropriately assess model training time in a live BSS context (as previously noted in [5]), one must consider not only model training time but also the total combined duration of EEG and EMG segments used in a single training run. In other words, the model training time is not merely the training length itself, but rather the maximum of the training length and the duration of neural data needed for a complete training run. In light of this, for the AT-AT system, we restricted training data to 120 EEG segments contaminated with high-variance EMG artifacts totaling 240s in combined duration. Thus, four minutes’ worth of data are required to support the AT-AT training runs described in this study.
@@
In accordance with methods for inferring signal SNR before running a subsequent tailored denoising model, we first run an initial upstream model based on a hybrid LSTM-CNN (LC) architecture to determine the appropriate SNR target level [5]; details of this upstream model are included in the Appendix. Subsequently, as a first denoising pass, the AT-AT architecture employs a convolutional autoencoder with batch normalization; this architecture showed promise for EEG-EMG denoising in prior work [5]. This autoencoder structure performs reasonably well in lower-noise settings but still demonstrates suboptimal performance on especially noisy reconstruction sites. Thus, to rectify the shortcomings of this convolutional autoencoder, we selectively invoke a time-series transformer model to handle more intensive instances of reconstruction. Introducing a second model architecture comes at the cost of a more complex parameter space search and longer training run, so the time-series transformer is only used when necessary in order to accommodate a more intensive retraining process. This novel hybrid use of autoencoders and transformers in conjunction is a key component of the AT-AT system; an evaluation of a sans-transformer method is outside of the scope of this study, but we refer readers to [5].
@@
One key challenge stemming from this combination of autoencoder and transformer architectures was that of spectral preservation. The selective insertion of time-series transformer-generated signals can introduce irregularities into the model output and prevent the model output characteristics from appropriately resembling the spectral characteristics of authentic EEG data. To achieve a well-regulated output compliant with the physical properties of EEG, we implemented an adversarial training component to our autoencoder-transformer hybrid architecture. Specifically, as mentioned in the Introduction, we treated the transformer-augmented autoencoder output as the “generator" and implemented a convolutional neural network-based (1D-CNN-based) discriminator model to distinguish between authentic versus model-reconstructed EEG. This generator-discriminator pair was folded into a five-cycle GAN training loop; ablation-wise, in line with prior work [5], we found that the addition of adversarial training led to $>$10% gains in spectral filtration accuracy.
After building and validating the AT-AT system across a partitioned 100-sample test set on the benchmark semi-synthetic contamination problem, we also implemented additional end-use case validation. In the EEGdenoiseNet context, performance can be assessed via direct comparison with the synthetically contaminated ground truth samples. But in real-world environments, the only way to assess the performance of our signal processing pipeline is by measuring improvement on downstream classification tasks. Thus, in order to demonstrate the potential of an AT-AT model at unlocking better BCI outcomes, we deployed AT-AT onto a notable BCI use case: digit- versus non-digit classification on the MindBigData EEG dataset [26] (see Appendix for further details). This downstream application ultimately incorporated an extensive geometric machine learning pipeline which is discussed in [3] (but falls outside of the scope of this paper).
@@ 3 Results & Discussion
Figure 3: AT-AT performance relative to major deep learning benchmarks. [A], [B], and [C] refer to [32, 30] and [8], respectively. All values are best extrapolated from reported documentation (best-in-class values in bold).
@@
Examining our system on real-world use cases is an additional key component of AT-AT validation. As discussed in the Appendix and formally treated in [3], we implemented AT-AT in conjunction with an extensive downstream manifold learning pipeline on the real-world imagined digit- versus non-digit classification problem. The results of this use case deployment (which are informally included in the Appendix) offer early evidence that AT-AT may confer utility in real-world BCI settings.